Suppose the marginal cost curve is given by the function q^2 + 5q + 30. If the demand curve is the same as before, what will be market price and market quantity?
1
Expert's answer
2021-07-20T09:59:11-0400

Suppose the demand curve is given by "P=120-.5Q"

"TR=P\times Q\\=(120-0.5Q)Q\\=120Q-0.5Q^2\\MR=120-Q"


"MC=Q^2+5Q+30"


"MR=MC\\120-Q=Q^2+5Q+30\\Q^2+5Q+Q+30-120=0\\Q^2+6Q-90=0"


To find Q we use the quadratic formula

"\frac{-b\pm\sqrt{b^2-4ac}}{2a}"


"\frac{-6\pm\sqrt{6^2-(4\times-90)}}{2}"


"-6\pm \frac{\sqrt{36+360}}{2}"


"\frac{-6\pm\sqrt{36+360}}{2}"


"\frac{-6\pm\sqrt{396}}{2}"


"\frac{-6+\sqrt{396}}{2}\space or\space \frac{-6-\sqrt{396}}{2}"


"=6.949874371\space or\space -12.94987437"


therefore "Q\approx7"

"P=120-.5(7)=116.5"